OpenStudy (anonymous):

Can someone explain to me how 6x^2(x + 4)^5 / 9x^3(x + 4) equals to 2(x + 4)^4 / 3x

Nnesha (nnesha):

\[\huge\rm \frac{ 6x^2(x+4)^5 }{ 9x^3(x+4) } = \frac{ 2(x+4)^4 }{ 3x }\] you can cancel out few things here hmm

OpenStudy (anonymous):

All I know is that 6 and 9 is 2/3 and x3 subtracts with x2, but i dont get how you get the 4th power

OpenStudy (anonymous):

and how the x+4 dont cancel out

Nnesha (nnesha):

\[\huge\rm \frac{ \color{red}{6x^2} \color{green}{(x+4)} \times (x+4)^4 }{ \color{red}{9x^3}\color{green}{(x+4)} } = \frac{ 2(x+4)^4 }{ 3x }\] you can cancel out few things here hmm

Nnesha (nnesha):

when you multiply same variables number you should add their exponent \[\huge\rm x^m + x^n = x^{m+n}\]

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @MUSTARD_MK All I know is that 6 and 9 is 2/3 and x3 subtracts with x2, but i dont get how you get the 4th power \(\color{blue}{\text{End of Quote}}\) which one ?? at the right side ?

OpenStudy (anonymous):

Yes sorry

Nnesha (nnesha):

nah that's fine it's in the question

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @MUSTARD_MK Can someone explain to me how 6x^2(x + 4)^5 / 9x^3(x + 4) equals to 2(x + 4)^4 / 3x \(\color{blue}{\text{End of Quote}}\) here equals to 2(x+4)^4 <----

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i dont get how it went from a 5 to a 4, that really confuses me

OpenStudy (anonymous):

I feel dumb lol

Nnesha (nnesha):

ohh left side ? like i said when you multiply same variables you should add their exponent like \[\huge\rm \frac{ x^2 }{ x^3 } = \frac{ \cancel x \times \cancel x }{ x \times \cancel x \times \cancel x }\]

Nnesha (nnesha):

so that's how (x+4)^5 can be written as (x+4) (x+4)^4 let's try by this way applying exponent rule \[\huge\rm \frac{ x^m }{ x^n } = x^{m -n}\] when bases are same you can move exponent from bottom to the top so \[\large\rm \frac{ (x+4)^5 }{ (x+4)^1 } = (x+4)^{5-1}\]

OpenStudy (anonymous):

Oooooooooh

OpenStudy (anonymous):

So theres like a hidden 1 there?

OpenStudy (anonymous):

sort of

Nnesha (nnesha):

yes thief 1 :/ :-) yes one is always there so if i ask what's the exponent of (x) answer would be ?

OpenStudy (anonymous):

1

Nnesha (nnesha):

- (x+ 5) how would you solve this distribute by what ? example

OpenStudy (anonymous):

-x -5 right?

Nnesha (nnesha):

yes so distribute by -1 right ? so yes 1 is always there :-)

OpenStudy (anonymous):

awesome

Nnesha (nnesha):

back to the question \[\huge\rm \frac{ \color{red}{6x^2} \color{green}{(x+4)} \times (x+4)^4 }{ \color{red}{9x^3}\color{green}{(x+4)} } = \frac{ 2(x+4)^4 }{ 3x }\]

OpenStudy (anonymous):

I don't want to piss you off, but where did the extra (x+4) from the top left came from?

Nnesha (nnesha):

hahha it's okay gO_Od to ask a question :-) i like that i just separated that

OpenStudy (anonymous):

Oooooh

OpenStudy (anonymous):

thats smart lol

Nnesha (nnesha):

okay we can write like this \[\huge\rm \frac{ \color{red}{6x^2} \times (x+4)^5 }{ \color{red}{9x^3}\color{green}{(x+4)^1} } = \frac{ 2(x+4)^4 }{ 3x }\] apply exponent rule

OpenStudy (anonymous):

ah, that makes a bit more sense now for me

Nnesha (nnesha):

yeah i thought you are familiar with exponent rules so it would help you to understand :-)

OpenStudy (anonymous):

Thanks man, I really appreciate it :)

Nnesha (nnesha):

got it ?? :o

Nnesha (nnesha):

finish ?? :o

OpenStudy (anonymous):

Yeah, I understand how to do it now

Nnesha (nnesha):

ohh okay nice ! gO_Od luck!!!

OpenStudy (anonymous):

Gracias!

Nnesha (nnesha):

if you want me to check ur answer just tag meh :-)

OpenStudy (anonymous):

I'll keep that noted

Nnesha (nnesha):

no hay problema